Joe can eat an Uranian sandwich in 25 minutes. Working together, Joe and Mark can eat the sandwich in 50/7 minutes. How many minutes would it take Mark to eat the sandwich by himself?
Answer by edjones(8007)

Let x=minutes it takes for Mark to eat a sandwich.
In 1 minute joe can eat 1/25 of a sandwich.
In 1 minute together they can eat 7/50 of a sandwich.
In 1 minute Mark can eat 1/x of a sandwich.
.
1/25 + 1/x = 7/50
2x+50=7x Multiply each side by 50x to eliminate fractions.
-5x=-50
x=10 min
.
Ed